OpenStudy (anonymous):

Which combination of integers can be used to generate the Pythagorean triple (5,12,13) ?

ganeshie8 (ganeshie8):

use the parametric formulas a = x^2 - y^2 b = 2xy c = x^2 + y^2

ganeshie8 (ganeshie8):

b = 2xy = 12 so xy = 6 since x>y, we get x=3, y=2

OpenStudy (anonymous):

thank you!!!

ganeshie8 (ganeshie8):

yw

OpenStudy (mathmath333):

https://www.youtube.com/watch?v=86YAPbZmsRI